OpenStudy (anonymous):

need the solutions √90 - √10

OpenStudy (anonymous):

Hi levie

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

2√10

OpenStudy (anonymous):

\[\sqrt{90} - \sqrt{10} = \sqrt{9(10)} -\sqrt{10} = \sqrt{9}\sqrt{10} - \sqrt{10} = 3\sqrt{10} - \sqrt{10} = 2\sqrt{10} \]

OpenStudy (anonymous):

lol, mental mistake

OpenStudy (anonymous):

I was trying to shortcut it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Thanks Joemath

OpenStudy (anonymous):

I was on my way to correcting the error

OpenStudy (anonymous):

joemath just didn't give me a chance, I have no room for error here

OpenStudy (anonymous):

thats ok Hero u help me a lots the other day.